OpenStudy (anonymous):

Please help! I give medals and become a fan! use de moivre's theorem to write the complex number in standard form. (sqrt2(cos(15)+i sin(15))^6

OpenStudy (anonymous):

a. 4-4sqrt3 i b. -8 c. 8i d. -4sqrt3 +4i

OpenStudy (psymon):

Well, when you want to find some power of a complex number in polar form: \[r(\cos \theta + isin \theta)^{n}\], we need to follow this formula: \[r ^{n}(\cos(n \theta) + isin(n \theta))\] SO basically, your power is 6. SO that means we must do: \[\sqrt{2}^{6}(\cos(6*15) + isin(6*15))\] Think you can get that for me and then we'll finish?

OpenStudy (anonymous):

8(cos(90)+isin(90))

OpenStudy (anonymous):

@Psymon what is the next step? :)

OpenStudy (psymon):

I apologize, doing too much work, haha. Okay, awesome. So now what is the cos(90) and sin(90)?